Month The following spreadsheet contains monthly returns for Cola Co. and Gas Co. for 2013. Using these data, estimate the avAlso, The volatility for Gas Co. is ___ % (Round to two decimal places.)

Thank you
Answer #1
Month Returns
coca cola (x) gas co.(y) x-a y-b (x-a)^2 (y-b)^2
jab 9.4 -3          7.46        -2.42                         55.61                         5.84
feb 2.2 -5.4          0.26        -4.82                           0.07                       23.20
mar -6.29 -4.1        -8.23        -3.52                         67.77                       12.37
aprl 7.4 0.4          5.46          0.98                         29.78                         0.97
may -0.1 -0.3        -2.04          0.28                           4.17                         0.08
jun -1.8 0.4        -3.74          0.98                         14.01                         0.97
jul 0.1 -2.5        -1.84        -1.92                           3.39                         3.67
aug 1.9 -2.1        -0.04        -1.52                           0.00                         2.30
sept 0.5 6.2        -1.44          6.78                           2.08                       46.01
oct -1.5 2.5        -3.44          3.08                         11.85                         9.51
nov 5.7 0.3          3.76          0.88                         14.12                         0.78
dec 5.8 0.6          3.86          1.18                         14.88                         1.40
total 23.31 -7                       217.74                    107.10
average monthly return 23.31/12 -7/12
                         1.94        -0.58
Let expected return be named as a b
Standard deviation/volatility [{(x-a)^2}/n]^(1/2) =(217.74/12)^(1/2) =(107.1/12)^(1/2)
                          4.26                         2.99
NOTE
n=no of months i.e.12
a=expected return of coca cola i.e.x
b=expected return of coca cola i.e.y

So volatility of gas co. is 2.99